OpenStudy (anonymous):

A rectangle is 6cm long and 5 cm wide. When each dimension is increased b x cm, the area is tripled. Fid the value of x.

OpenStudy (anonymous):

(x+6)(x+5) = 3(30) Solve.

OpenStudy (anonymous):

4

OpenStudy (anonymous):

add four to 6 and four to 5: 10(9)=90 so four is the right answer, I solved the equation. :)

OpenStudy (anonymous):

x^2 +11x +30=90 ?